# KeyB2: Selecting Key Blocks is Also Important for Long Document Ranking with Large Language Models

### v1

@@ 5.4. Fine-tuning on TREC DL19 Document Ranking Task
The findings suggest that combining LLMs with a key block selection strategy largely enhances retrieval performance for long document ranking tasks. This answers RQ3. The improvements with cross-encoder configurations indicate that leveraging semantic-rich representations for relevance matching is more effective, partly answers RQ2. The consistent gains across different KeyB2 settings confirm the robustness and adaptability of the proposed approach.
In terms of Llama2 and Llama3, on this dataset, it shows similar results.
5.5. Zero-Shot Performance on Robust04
We load the checkpoint of each model trained in Section 5.4 and evaluate each model on Robust04.
Results: As shown in Table 3, the previous KeyB(BERT)BM25 model performs worse than BM25 on average, whereas the KeyB2 models maintain competitive performance without additional training on Robust04. The KeyB2 models all demonstrate superior zero-shot performance compared to RankLLaMA. Notably, KeyB2(Llama3)cross achieved the highest scores, with NDCG@20 of 0.4887 and P@20 of 0.4187, followed closely by KeyB2(Llama3)BM25, which reached NDCG@20 of 0.4758 and P@20 of 0.4129. In comparison, RankLLaMA recorded NDCG@20 of 0.4472 and P@20 of 0.3873. These results indicate that the KeyB2 models, especially the Llama3-based configurations, generalize effectively to unseen datasets.
Significance Testing: We perform significant tests on the Robust04 dataset, comparing KeyB2 models with RankLLaMA and previous BERT-based KeyB models. The results indicate that KeyB2 models using the cross-encoder block selection strategy significantly outperformed RankLLaMA across all metrics. Over half of the metrics for BM25 or bi-encoder strategies also showed significant improvement over RankLLaMA. However, when compared to KeyB(BERT)BinB, only the KeyB2(Llama3)cross configuration displayed significant improvements across most metrics. When comparing KeyB2 models using Llama3 with using Llama2, only KeyB2(Llama3)cross show significant differences as well.
The zero-shot evaluation results demonstrate that KeyB2 models, particularly those utilizing the Llama3 model with cross-encoder strategies obtains the best results. While the significant tests of other KeyB2 models, suggest further investigation might be needed. This leads to the question of whether fine-tuning on Robust04 might yield better findings.
Table 3. Zeroshot Results on Robust04 dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 P@5 P@10 P@20 MAP NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG
Baseline models
BM25 0.5542 0.5004 0.4382 0.3631 0.2334 0.5080 0.4741 0.4485 0.4240 0.4402
BERT based models
KeyB(BERT)BM25 0.5502 0.4787 0.4368 0.3716 0.2226 0.5080 0.4652 0.4461 0.4258 0.4196
KeyB(BERT)BinB 0.6107 0.5269 0.4699 0.3958 0.2444 0.5705 0.5087 0.4845 0.4606 0.4369
LLM
RankLLaMA 0.5663 0.5100 0.4518 0.3873 0.2296 0.5241 0.4904 0.4646 0.4472 0.4269
KeyB2 models
KeyB2(Llama2)BM25 0.5662 0.5357R 0.4880R 0.4074R 0.2404R 0.5341 0.5161R 0.4981R 0.4713R 0.4356R
KeyB2(Llama2)cross 0.6265R 0.5349R 0.4739R 0.4020R 0.2406R 0.5843R 0.5241R 0.4934R 0.4692R 0.4378R
KeyB2(Llama2)bi 0.5984 0.5422R 0.4723R 0.3994R 0.2384R 0.5522 0.5228R 0.4899R 0.4655R 0.4347R
KeyB2(Llama3)BM25 0.6024 0.5390R 0.4888R 0.4129RB 0.2472R† 0.5602 0.5191R 0.5006R 0.4758R 0.4390R
KeyB2(Llama3)cross 0.6305R 0.5622RB† 0.4996RB† 0.4187RB† 0.2520R† 0.5904R 0.5419RB 0.5162RB† 0.4887RB† 0.4462RB†
KeyB2(Llama3)bi 0.5743 0.5213 0.4779R 0.4074R 0.2437R 0.5341 0.5043 0.4895R 0.4693R 0.4381R
5.6. Fine-tuning on Robust04 and MLDR-zh
We now want to compare the results of KeyB2 models after fine-tuning on Robust04, not the zero-shot setting.
Table 4. Fine-tuned models’ results on Robust04 dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 P@5 P@10 P@20 MAP NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG
Baseline models
BM25 0.5542 0.5004 0.4382 0.3631 0.2334 0.5080 0.4741 0.4485 0.4240 0.4402
BERT based models
KeyB(BERT)BM25 0.6347 0.5445 0.4887 0.4161 0.2473 0.5924 0.5311 0.5043 0.4805 0.4404
KeyB(BERT)BinB 0.6625 0.5614 0.5007 0.4226 0.2556 0.6243 0.5446 0.5163 0.4902 0.4458
LLM
RankLLaMA 0.5700 0.5412 0.4947 0.4218 0.2508 0.5359 0.5178 0.5005 0.4804 0.4377
KeyB2 models
KeyB2(Llama2)BM25 0.6744R 0.6063RB 0.5272R 0.4447RB 0.2668R 0.6222R 0.5799RB 0.5416R 0.5140RB 0.4521R
KeyB2(Llama2)cross 0.6586R 0.5758R 0.5023 0.4329 0.2554 0.6024R 0.5546R 0.5201 0.4992R 0.4464R
KeyB2(Llama2)bi 0.6504R 0.5959R 0.5348RB 0.4482RB 0.2691R 0.6023R 0.5728R 0.5458RB 0.5168RB 0.4553R
KeyB2(Llama3)BM25 0.7186R 0.6289RB 0.5643RB† 0.4749RB† 0.2822RB† 0.6684R 0.6033RB 0.5755RB† 0.5472RB† 0.4654RB†
KeyB2(Llama3)cross 0.7313R† 0.6514RB† 0.5811RB† 0.4799RB† 0.2923RB† 0.6851R† 0.6302RB† 0.5977RB† 0.5593RB† 0.4718RB†
KeyB2(Llama3)bi 0.6984R 0.6264RB† 0.5665RB† 0.4764RB† 0.2883RB† 0.6502RB 0.6006RB 0.5748RB† 0.5461RB† 0.4654RB†
Results on Robust04: As shown in Table 4, all neural models largely outperform the BM25 results. RankLLaMA obtains lower results than KeyB(BERT)BinB, indicating directly inputting a long document to LLMs might not be a good solution. The proposed KeyB2(Llama3)cross model achieve the best performance on all metrics, with NDCG@20 = 0.5593 and P@20 = 0.4799, surpassing all baseline models, including RankLLaMA with NDCG@20 = 0.4804 and P@20 = 0.4218, the best KeyB model KeyB(BERT)BinB with NDCG@20 = 0.4902 and P@20 = 0.4226. In terms different Llama base models, the KeyB2 using Llama3 obtains overall better results.
Results on MLDR-zh: As shown in Table 5, all models outperform the BM25 baseline by a significant margin. RankLLaMA scores are lower than those of both KeyB models with BERT, suggesting that directly feeding long documents to LLMs may not be the most effective approach for this task. The KeyB2(Llama2)bi configuration obtains the best results across all metrics, with NDCG@8 = 0.9744 and MAP = 0.9661, outperforming all baseline models. The MAP score is higher than RankLLaMA of 115.4%, whose is 0.4485, and higher than the good-performing KeyB(BERT)BinB model of 10.3%, which reaches 0.8758. On MLDR-zh dataset, the bi-encoder strategy looks overall better.
Significance Testing: On fine-tuned Robust04, comparing with RankLLaMA, all models significantly outperform it, with most metrics of KeyB2(Llama2)cross and all metrics of other KeyB2 models. Comparing with the best KeyB model KeyB(BERT)BinB, four metrics of KeyB2(Llama2)BM25 and KeyB2(Llama2)bi significantly improve it, and most metrics of KeyB2(Llama3) models show significant differences. In terms of using Llama2 or Llama3, all metrics of cross-encoder strategy shows significant difference and the other two shows significant improvements with 6 metrics.
On MLDR-zh dataset, all metriecs of proposed KeyB2 models show significant improvements comparing with RankLLaMA and KeyB(BERT)BinB. In terms of comparing models using Llama2 and Llama3, KeyB2 models with BM25 or cross-encoder show significant differences.
Key Observations: On these two datasets, the fine-tuned results show the proposed KeyB2 models can significantly improve the baseline models, including RankLLaMA which directly input long sequences. These results address RQ3, demonstrating the importance of key block selection in the era of LLMs. In terms of different strategies, results show either using cross-encoder or bi-encoder would obtain the best results, on different datasets. All strategies can be successfully incorporated in KeyB2 models, showing the flexibility and adaptability. Comparing with the zero-shot performance, fine-tuning on Robust04 shows larger improvements, and better potential of using newer Llama model.
Table 5. Results on MLDR Chinese dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 MAP NDCG@1 NDCG@5 NDCG@8
Baseline models
BM25 0.2013 0.2591 0.2013 0.2629 0.2768
BERT based models
KeyB(BERT)BM25 0.4225 0.5862 0.4225 0.6249 0.6846
KeyB(BERT)BinB 0.8038 0.8758 0.8038 0.9001 0.9066
LLM
RankLLaMA 0.2975 0.4485 0.2975 0.4439 0.5752
KeyB2 models
KeyB2(Llama2)BM25 0.8638RB 0.9099RB 0.8638RB 0.9231RB 0.9319RB
KeyB2(Llama2)cross 0.9038RB 0.9394RB 0.9038RB 0.9497RB 0.9543RB
KeyB2(Llama2)bi 0.9475RB 0.9661RB 0.9475RB 0.9706RB 0.9744RB
KeyB2(Llama3)BM25 0.9088RB† 0.9420RB† 0.9088RB† 0.9520RB† 0.9563RB†
KeyB2(Llama3)cross 0.9325RB† 0.9584RB† 0.9325RB† 0.9653RB† 0.9688RB†
KeyB2(Llama3)bi 0.9462RB 0.9648RB 0.9462RB 0.9705RB 0.9735RB
@@ 5.7. Analysis of the Position of Selected Blocks
Figure 6. The probabilities of top 8 block appearing locations of TREC 2019 DL with different block selecting methods.
Figure 7. The probabilities of top 8 block appearing locations of Robust04 with different block selecting methods.
Figure 8. The probabilities of top 8 block appearing locations of MLDR-zh with different block selecting methods.
@@
We analyze the block selection patterns across different strategies (BM25, cross-encoder, and bi-encoder) and datasets, as shown in Fig. 6, Fig. 7, and Fig. 8. These figures show the probability distributions of the top 8 selected blocks across various document positions, highlighting trends in how different methods prioritize blocks within long documents.
Dataset-Specific Observations
@@
  • TREC DL19 (Fig. 6): All three methods show similar preferences for blocks in the front half. The cross-encoder focuses more on the front half, while the BM25 approach continues to emphasize the later positions, particularly the last three.
@@
  • Robust04 (Fig. 7): While early positions are still preferred, middle and later parts contribute more, particularly with the cross-encoder and bi-encoder methods. These strategies show a more distributed selection, retrieving relevant information from various positions and displaying a more spread-out pattern compared to TREC DL19.
@@
  • MLDR-zh (Fig. 8): All three methods exhibit a stronger focus on early positions, with fewer blocks selected from later parts. This trend suggests that, for this Chinese dataset, relevant information may be more concentrated in the front positions. However, relevant information are still found and used in other positions, also contributing large portions.
General Trends

### v2

@@ 6.2. Cross-Dataset Effectiveness Summary (Answer to RQ2)
  • Significant wins on every corpus: The winning KeyB2 variant is statistically superior to both baselines on all five evaluation settings, as indicated by the double superscript R,B.
[…]
  • Successful design: Comparing with RankLLaMA-MaxP and AvgP, the score pooling models, which perform relatively poor, the proposed select-then-combine-then-rerank design strategy of KeyB2 shows large improvements.
  • Generalisation: Without in-domain tuning, KeyB2 still delivers significant gains on Robust04 (+8.9% NDCG@10); after fine-tuning the advantage widens to +25.4%.
  • Language robustness: The framework remains highly effective on MLDR-zh where documents average over 12k tokens, confirming that KeyB2 is not limited to English data.
Conclusion for RQ2: Across five diverse evaluation settings, and with statistical significance in every case, KeyB2 which pairs an LLM reranker with an explicit block selector consistently outperforms both the full-document LLM baseline RankLLaMA (including MaxP, AvgP) and the strongest BERT-based block-selection system KeyB(BERT)BinB, establishing the successful design (select-then-combine-then-rerank) of KeyB2, as a clearly more effective strategy for long-document ranking.
6.3. Efficiency and Resource Consumption
We compare the reranking time and GPU memory requirements of different models, as shown in Fig. 8. Each experiment is measured on a server with a NVIDIA A100 40GB GPU, Intel Xeon 6248R CPU, and 128GB memory, using the TREC 2019 DL document reranking dataset. Fig. 8(a) shows the average reranking time for 100 documents per query (note that the bi-encoder version precomputes block embeddings offline for fast online search). Fig. 8(b) records the GPU memory usage during efficient LoRA fine-tuning without gradient checkpointing with batch size 1. This setup demonstrates the minimum GPU requirements for training without advanced techniques like DeepSpeed ZeRO (Rasley et al., 2020; Rajbhandari et al., 2020) or gradient checkpointing, which may require multiple GPUs or reduce training efficiency.
(a) Avg. time (seconds) to rerank 100 documents given a query.
(b) Peak GPU memory during training (MB). Note: RankLLaMA with batch size 1 exceeds 96GB memory (H20 GPU).
Figure 8. Efficiency comparisons on TREC 2019 DL.
6.3.1. Inference Time Comparison
The full-document RankLLaMA model requires 50.65 seconds to rerank 100 documents, making it the slowest approach due to autoregressive decoding over long inputs. In contrast, KeyB2 models significantly reduce inference latency by pre-selecting relevant blocks:
  • KeyB2(Llama2)bi: 6.25 seconds (87.7% faster than RankLLaMA)
  • KeyB2(Llama3)cross: 8.58 seconds (83.1% faster)
  • KeyB2-BM25 variants: ¡6 seconds (fastest, over 88% faster)
For comparison, RankLLaMA-MaxP (11.6 seconds) and AvgP (11.5 seconds), which average over segmented blocks, offer moderate efficiency improvements than RankLLaMA but remain around 35% slower than the slowest KeyB2 variant. These results highlight that block selection reduces inference time more effectively than post-hoc score aggregation.
6.3.2. Memory Requirements
Running RankLLaMA on full 4096-token documents requires over 96GB of GPU memory even with batch size 1, making training infeasible on standard GPUs without aggressive techniques such as gradient checkpointing, which introduce substantial overhead.
KeyB2 models drastically reduce memory demands:
  • KeyB2(Llama2)BM25: 22,523 MB (over 77% reduction vs. RankLLaMA)
  • KeyB2(Llama3)cross: 25,297 MB (over 74% reduction vs. RankLLaMA)
RankLLaMA-MaxP and AvgP are the most memory-efficient (17,235 MB), but as shown in Section 6.1, this comes at the cost of notable performance drops. KeyB2 models, by contrast, retain high accuracy while fitting comfortably on widely available hardware (e.g., NVIDIA 4090 for Llama2, V100/5090 for Llama3).
6.3.3. Pareto Frontier
Figure 9 plots NDCG@10 against latency and memory. All KeyB2 variants form or dominate the Pareto frontier in terms of speed: BM25 sits at the extreme efficiency corner, the bi-encoder offers the best quality–cost trade-off on the 7B backbone, and the cross-encoder pushes effectiveness to the maximum on 8B with acceptable overhead.
Conclusion for RQ4: Block selection drastically reduces computational overhead relative to full-document LLM reranker, RankLLaMA:
[…]
  • 74–77 % lower GPU memory during standard LoRA fine-tuning (>96>\!96> 96 GB $\to$ 22–25 GB) [^14] Without gradient checkpointing, which would slow training considerably..
Compared to RankLLaMA, which lies far outside the Pareto frontier, all KeyB2 variants offer superior accuracy-efficiency trade-offs. They consistently reduce computational cost while improving ranking effectiveness, thereby establishing a new Pareto frontier for LLM-based long-document reranking on a single 24–32 GB GPU.
(a) NDCG@10 vs. reranking time (100 docs).
(b) NDCG@10 vs. training memory (batch size 1).
Figure 9. Accuracy-efficiency trade-off on TREC 2019 DL (top-left is optimal).
Here we give an explanation about the above efficiency results. Although recent LLMs can handle input sizes from thousands to hundreds of thousands of tokens, their computational complexity still depends on the input token count. Similar to the Transformer or BERT architectures, recent LLMs like Llama2 or Llama3 also use self-attention mechanism. The difference is recent LLMs use the decoder-only architecture with unidirectional attention. Llama models use Grouped Query Attention (GQA) (Ainslie et al., 2023): instead of each attention head having its own key (K) and value (V) matrices, multiple heads are grouped together, and each group shares the same key and value vectors. This mechanism is designed to reduce memory and computational overhead, particularly in large models like Llama and Mistral (Jiang et al., 2023), by allowing several query heads to reuse the same key-value pair within a group. This helps speed up inference while maintaining the benefits of multi-head attention. Recent LLMs also use KV Cache (Pope et al., 2023) to speed up inference, but the KV Cache still faces memory bound issue. Besides, the cache does not apply to RankLLaMA. Indeed, the KV Cache technique caches the Key (K) and Value (V) matrices from previous tokens, enabling the model to reuse these stored results when generating new tokens. This avoids redundant computations and significantly improves efficiency. However, the RankLLaMA model only use the last token to infer the relevance once and does not need to generate more tokens. Based on the above analyses, the complexity of the RankLLaMA model remains quadratic relative to the input token count. On the contrary, KeyB2 models use addtional small models to filter out many irrelevant tokens, resulting quadratic efficiency improvement.
6.4. Selector Design and Effectiveness–Efficiency Trade‐off (RQ3)
Combining the effectiveness tables (Section 6.1) with the efficiency profiles (Section 6.3) yields the observations:
  1. (1)
    On TREC DL datasets, the cross-encoder is the strongest selector. The cross-encoder selector overall outperforms both BM25 and the bi-encoder on the TREC DL19 and TREC DL23 datasets. Specifically, on TREC DL19, KeyB2(Llama3)cross significantly outperforms KeyB2(Llama3)bm25, while KeyB2(Llama3)bi shows no significant difference from BM25. On TREC DL23, KeyB2(Llama2)cross also performs better and is significantly better than KeyB2(Llama2)BM25, whereas KeyB2(Llama2)bi is not.
    In addition, across datasets, the cross-encoder variant demonstrates the highest performance potential and peak results (e.g., on TREC DL and Robust04). This improvement comes at the cost of approximately 29.3% to 45.0% increased latency and only around 90 MB additional memory usage than other selectors.
  2. (2)
    On the Robust04 dataset, BM25 can be relatively strong. Under the zero-shot and Llama2 setting, KeyB2(Llama2)BM25 achieves the best performance, and KeyB2(Llama3)BM25 performs better than KeyB2(Llama3)bi.
    Under the fine-tuned setting, KeyB2(Llama2)BM25 performs better than KeyB2(Llama2)cross, and KeyB2(Llama3)BM25 performs similarly to KeyB2(Llama3)bi.
    This may be because Robust04 documents average just over 700 tokens in length (see Table 2). As a result, the benefit of sophisticated learned selectors is limited. Block-level selection plays a less critical role, and BM25’s lexical signals may suffice to identify relevant content. This highlights a scenario where BM25-based selection remains competitive, especially when document length is moderate.
  3. (3)
    On the MLDR-zh dataset, bi-encoders generally perform better. KeyB2(Llama2)bi shows significant improvements over both KeyB2(Llama2)bm25 and KeyB2(Llama2)cross. KeyB2(Llama3)bi performs similarly to the cross-encoder one, and both are significantly better than KeyB2(Llama3)BM25.
    This indicates that bi-encoders can be particularly effective in certain settings, especially when the model is well-trained on the target language or domain, and can already capture sufficient semantic relevance. In such cases, the advantage of cross-encoders may not manifest, highlighting the importance of language and domain alignment in selector performance.
  4. (4)
    Efficiency and deployment trade-offs among selectors. In terms of inference efficiency, BM25 is the fastest and most lightweight option. It requires no model parameters or preprocessing, can be computed directly online using inverted indexes, and integrates seamlessly with KeyB2’s pipeline.
    Bi-encoder selectors offer a good balance between effectiveness and latency. Once block embeddings are pre-computed, online selection becomes fast and memory-efficient. However, this approach introduces engineering complexity and storage overhead due to the need for maintaining an external embedding index.
    Cross-encoders, while achieving the best accuracy on most English datasets, incur the highest runtime cost. They require scoring each block on-the-fly with the LLM, resulting in approximately 39–45% longer latency than BM25. Nevertheless, they are plug-and-play, requiring no offline indexing or embedding preparation, and incur only minimal memory overhead (around 90 MB compared to bi-encoder or BM25).
Answer to RQ3: Selector Design. Given the observations, we find that no single block selector dominates universally; instead, each offers a different balance of accuracy, efficiency, and engineering cost.
Cross-encoder deliver the highest accuracy across most English benchmarks, such as TREC DL19 and DL23, and achieve the strongest performance peak overall. It excels when document length exceeds (e.g., DL19 $\sim$1.9k tokens) and token-level relevance modeling is needed. However, this comes at the cost of 29%–45% additional latency than bi-encoder and BM25, requiring real-time scoring of all blocks.
Bi-encoder offers a competitive alternative when the model is pretrained with strong alignment, as shown in MLDR-zh. It is more efficient once block embeddings are precomputed, but demand additional storage and indexing efforts. Its effectiveness varies across tasks—while strong on MLDR-zh, it show less consistent gains on TREC DL or Robust04.
BM25 remains a surprisingly strong baseline under certain conditions. On short documents (e.g., Robust04, $\sim$700 tokens), where block selection is less critical, BM25 performs comparably or even better than learned selectors. It requires no training or offline computation and is the most deployment-friendly choice, particularly for latency-sensitive systems.
Overall, the block selector in KeyB2 serves as a tunable component to navigate the accuracy–efficiency spectrum:
  • Use a cross-encoder when accuracy is the top priority and infrastructure supports modest latency increase.
  • Use a bi-encoder when seeking a balance of speed and effectiveness, particularly when it is good enough under some settings.
  • Use BM25 for maximum efficiency and simplicity.
These findings reinforce that selector design should be co-tuned with the LLM’s capacity, selector’s capacity, document length, and deployment constraints. KeyB2 enables flexible adaptation to such needs without compromising overall ranking quality.
@@ 6.5. Analysis of the Position of Selected Blocks
(a) TREC 2019 DL (MS MARCO).
(b) TREC 2023 DL (MS MARCO v2).
(c) Robust04.
(d) MLDR-zh.
@@
Figure 10. The probabilities of top-8 block appearing locations across different datasets with various block selecting methods.
@@
We analyze the block selection patterns across different strategies (BM25, cross-encoder, and bi-encoder) and datasets, as shown in Fig. 10. These figures show the probability distributions of the top 8 selected blocks across various document positions, highlighting trends in how different methods prioritize blocks within long documents.
6.5.1. Dataset-Specific Observations
@@
  • TREC DL19 (MS MARCO) (Fig. 10(a)): All three methods show similar preferences for blocks in the front half. The cross-encoder focuses more on the front half, while the BM25 approach continues to emphasize the later positions, particularly the last three.
  • TREC DL23 (MS MARCO v2) (Fig. 10(b)): Similar patterns are observed as in DL19, where all three methods predominantly select blocks from the first half. Compared to MS MARCO v1, MS MARCO v2 shows an even stronger concentration of selection in the early blocks, suggesting that relevant information in MS MARCO v2 tends to appear earlier within documents, while later blocks likely contain more non-relevant or noisy content.
@@
  • Robust04 (Fig. 10(c)): While early positions are still preferred, middle and later parts contribute more, particularly with the cross-encoder and bi-encoder methods. These strategies show a more distributed selection, retrieving relevant information from various positions and displaying a more spread-out pattern compared to TREC DL19.
@@
  • MLDR-zh (Fig. 10(d)): All three methods exhibit a stronger focus on early positions, with fewer blocks selected from later parts. This trend suggests that, for this Chinese dataset, relevant information may be more concentrated in the front positions. However, relevant information are still found and used in other positions, also contributing large portions.
6.5.2. General Trends